# The Landscape of Agentic Reinforcement Learning for LLMs: A Survey

### v1

@@ Abstract
Abstract: The emergence of agentic reinforcement learning (Agentic RL) marks a paradigm shift from conventional reinforcement learning applied to large language models (LLM RL), reframing LLMs from passive sequence generators into autonomous, decision-making agents embedded in complex, dynamic worlds. This survey formalizes this conceptual shift by contrasting the degenerate single-step Markov Decision Processes (MDPs) of LLM-RL with the partially observable, temporally extended partially observable Markov decision process (POMDP) that define Agentic RL. Building on this foundation, we propose a comprehensive twofold taxonomy: one organized around core agentic capabilities, including planning, tool use, memory, reasoning, self-improvement, and perception, and the other around their applications across diverse task domains. Central to our thesis is that reinforcement learning serves as the critical mechanism for transforming these capabilities from static, heuristic modules into adaptive, robust agentic behavior. To support and accelerate future research, we consolidate the landscape of open-source environments, benchmarks, and frameworks into a practical compendium. By synthesizing over five hundred recent works, this survey charts the contours of this rapidly evolving field and highlights the opportunities and challenges that will shape the development of scalable, general-purpose AI agents.
Keywords: Agentic Reinforcement Learning, Large Language Models, LLM Agent
= Date: September 1st, 2025
[Uncaptioned image] Code Repository: https://github.com/xhyumiracle/Awesome-AgenticLLM-RL-Papers
= Corresponding: jeremyyin@robots.ox.ac.uk, bailei@pjlab.org.cn
= Main Contact: guibinz@u.nus.edu, genghejia0530@gmail.com, x.yu21@imperial.ac.uk
@@ 1 Introduction
[…]
my agentic tree/.style= forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=4pt, line width=1.1pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=10.5em,font=,, where level=2text width=11.5em,font=,, where level=3text width=12em,font=,, where level=4text width=50em,font=,,
The rapid convergence of large language models (LLMs) and reinforcement learning (RL) has precipitated a fundamental transformation in how language models are conceived, trained, and deployed. Early LLM-RL paradigms largely treated these models as static conditional generators, optimized to produce single-turn outputs aligned with human preferences or benchmark scores. While successful for alignment and instruction following, such approaches overlook the broader spectrum of sequential decision-making that underpins realistic, interactive settings. These limitations have prompted a shift in perspective: rather than viewing LLMs as passive text emitters, recent developments increasingly frame them as agentic entities, i.e.,, autonomous decision-makers capable of perceiving, reasoning, planning, invoking tools, maintaining memory, and adapting strategies over extended horizons in partially observable, dynamic environments. We define this emerging paradigm as Agentic Reinforcement Learning (Agentic RL). To more clearly delineate the distinction between the concept of Agentic RL studied in this work and conventional RL approaches, we provide the following definition:
@@
In Section 2, we present a more formal, symbolically grounded distinction between agentic RL and conventional RL. Prior research relevant to agentic RL can be broadly grouped into two complementary threads: LLM agents and RL for LLMs, detailed as follows:
@@
LLM-based agents represent an emerging paradigm in which LLMs act as autonomous or semi-autonomous decision-making entities [1, 2], capable of reasoning, planning, and executing actions in pursuit of complex goals. Recent surveys have sought to map this landscape from complementary perspectives. Luo et al. [3] propose a methodology-centered taxonomy that connects architectural foundations, collaboration mechanisms, and evolutionary pathways, while Plaat et al. [4] emphasize the core capabilities of reasoning, acting, and interacting as defining features of agentic LLMs. Tool use, encompassing retrieval-augmented generation (RAG) and API utilization, is a central paradigm, extensively discussed in Li et al. [5] and further conceptualized by Wang et al. [6]. Planning and reasoning strategies form another pillar, with surveys such as Masterman et al. [7] highlighting common design patterns like plan-execute-reflect loops, while Tao et al. [8] extend this to self-evolution, where agents iteratively refine knowledge and strategies without substantial human intervention. Other directions explore collaborative, cross-modal, and embodied settings, from multi-agent systems [9] to multimodal integration [10], and brain-inspired architectures with memory and perception [11].
RL for LLMs.
The second line of research investigates how reinforcement learning algorithms are applied to improve or align LLMs. A primary branch, RL for LLMs, leverages on-policy (e.g., proximal policy optimization (PPO) [12] and Group Relative Policy Optimization (GRPO) [13]) and off-policy (e.g., actor–critic, Q-learning [14]) methods to enhance capabilities such as instruction following, ethical alignment, and code generation [15, 16, 17]. A complementary direction, LLM for RL, examines the deployment of LLMs as planners, reward designers, goal generators, or information processors to improve sample efficiency, generalization, and multi-task planning in control environments, with systematic taxonomies provided by Cao et al. [18]. RL has also been integrated throughout the LLM lifecycle: from data generation [19, 20] and pretraining [21] to post-training and inference [22], as surveyed by Guo et al. [23]. The most prominent branch here is post-training alignment, notably Reinforcement Learning from Human Feedback (RLHF) [24], along with extensions such as Reinforcement Learning from AI Feedback (RLAIF) and Direct Preference Optimization (DPO) [25, 26, 27, 15].
@@
The recent surge in research on LLM agents and RL-enhanced LLMs reflects two complementary perspectives: one explores what large language models can do as the core of autonomous agents, while the other focuses on how reinforcement learning can optimize their behavior. However, despite the breadth of existing work, a unified treatment of agentic RL, which conceptualizes LLMs as policy-optimized agents embedded in sequential decision processes, remains lacking. Current studies often examine isolated capabilities, domains, or custom environments, with inconsistent terminology and evaluation protocols, making systematic comparison and cross-domain generalization difficult. To bridge this gap, we present a coherent synthesis that connects theoretical foundations with algorithmic approaches and practical systems. We formalize agentic RL through Markov decision process (MDP) and partially observable Markov decision process (POMDP) abstractions to distinguish it from classical LLM-RL paradigms, and introduce a capability-centered taxonomy that includes planning, tool use, memory, reasoning, reflection (self-improvement), and interaction as RL-optimizable components. Furthermore, we consolidate representative tasks, environments, frameworks, and benchmarks that support agentic LLM training and evaluation, and conclude by discussing open challenges and outlining promising future directions for scalable, general-purpose agentic intelligence. Overall, we aim to further clarify the research scope of this survey:
Primary focus: how RL empowers LLM-based agents (or, LLMs with agentic characteristics) in dynamic environments Out of scope (though occasionally mentioned): RL for human value alignment (e.g., RL for harmful query refusal); traditional RL algorithms that are not LLM-based (e.g., MARL [28]); RL for boosting pure LLM performance on static benchmarks.
@@
This survey is organized to progressively build a unified understanding of Agentic RL from conceptual foundations to practical implementations. The following structure is as follows: Section 2 formalizes the paradigm shift to Agentic RL through an MDP/POMDP lens. Section 3 examines agentic RL from the capability perspective, categorizing key modules such as planning, reasoning, tool using, memory, self-improvement, perception, and others. Section 6 explores applications across domains, including search, GUI navigation, code generation, mathematical reasoning, and multi-agent systems. Section 5 consolidates open-source environments and RL frameworks that underpin experimentation and benchmarking. Section 6 discusses open challenges and future directions towards scalable, adaptive, and reliable agentic intelligence, and Section 7 concludes the survey. The overall structure is also illustrated in Figure 1.
[…]
@@ 2 Preliminary: From LLM RL to Agentic RL
LLMs are initially pre-trained using behavior cloning, which applies maximum likelihood estimation (MLE) to static datasets such as web-scraped text corpora. Subsequent post-training methods enhance capabilities and align outputs with human preferences—transforming them beyond generic web-data replicators. A common technique is supervised fine-tuning (SFT), where models are refined on human-generated (prompt, response) demonstrations. However, procuring sufficient high-quality SFT data remains challenging. Reinforcement fine-tuning (RFT) offers an alternative by optimizing models through reward functions, circumventing dependence on behavioral demonstrations.
In early RFT research, the core objective is to optimize LLMs through human feedback [24] or data preferences [29], aligning them with human preferences or directly with data preferences (as in DPO). This preference-based RFT (PBRFT) primarily involves learning reward model optimization for LLMs on a fixed preference dataset, or directly implementing it using data preferences. With the release of LLMs such as OpenAI o1 [30] and DeepSeek-R1 [31] that possess reasoning capabilities, their improved performance and cross-domain generalization have garnered widespread attention. With the release of models like OpenAI o3 [32], which possess both self-evolving reasoning capabilities and support for tool use, researchers are beginning to contemplate how to deeply integrate LLMs with downstream tasks through reinforcement learning methods. Subsequently, researchers have shifted their focus from PBRFT, aimed at optimizing fixed preference datasets, to agentic reinforcement learning tailored for specific tasks and dynamic environments.
In this section, we provide a formalization of the paradigm shift from PBRFT to the emerging framework of agentic reinforcement learning (Agentic RL). While both approaches leverage RL techniques to improve LLMs’ performance, they fundamentally differ in their underlying assumptions, task structure, and decision-making granularity. Figure [33] illustrates the paradigm shift from LLM RL to agentic RL.
Figure 2: Paradigm shift from LLM-RL to agentic RL.
@@ 2.1 Markov Decision Processes
The Markov decision process (MDP) for the RL fine-tuning process can be formalized as a seven-element tuple $\langle\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{P},\mathcal{R},T,\gamma\rangle$, where $\mathcal{S}$ represents the state space and $\mathcal{O}$ is the observation space of the agent. $\mathcal{A}$ denotes the action space. $\mathcal{R}$ is defined as the reward function, $\mathcal{P}$ encapsulates the state transition probabilities, $T$ signifies the task horizon, and $\gamma$ is the discount factor. By casting both preference-based RFT and agentic RL as MDP or POMDP, we clarify the theoretical implications of treating LLMs either as static sequence generators or as interactive, decision-capable agents embedded within dynamic environments.
@@
$\langle\mathcal{S}_{\text{trad}},\mathcal{A}_{\text{trad}},\mathcal{P}_{\text{trad}},\mathcal{R}_{\text{trad}},T=1\rangle.$ (1)
@@
$\langle\mathcal{S}_{\text{agent}},\mathcal{A}_{\text{agent}},\mathcal{P}_{\text{agent}},\mathcal{R}_{\text{agent}},\gamma,\mathcal{O}\rangle,$ (2)
@@
$\mathcal{A}$: Action space Pure text sequence. $\mathcal{A}_{\text{text}}\cup\mathcal{A}_{\text{action}}$.
$\mathcal{P}$: Transition Deterministic to the terminal state. Dynamic transition function $P(s_{t+1}\mid s_{t},a_{t})$.
$\mathcal{R}$: Reward Single scalar $r(a)$. Step-wise $R(s_{t},a_{t})$; combines sparse task and dense sub-rewards.
@@ 2.2 Environment State
The LLM agent acts over multiple time-steps in a POMDP. Let $s_{t}\in\mathcal{S}_{\text{agent}}$ denote the full world-state and the LLM agent gets observation $O_{t}$ based on current state $o_{t}=\mathcal{O}(s_{t})$. The LLM agent chooses an action $a_{t}$ based on the current observation $o_{t}$, and the state evolves over time:
$s_{t+1}\sim P(s_{t+1}\mid s_{t},a_{t}),$ (4)
@@ 2.3 Action Space
In the agentic RL setting, the LLM’s action space comprises two disjoint components:
$\mathcal{A}_{\text{agent}}=\mathcal{A}_{\text{text}}\cup\mathcal{A}_{\text{action}}$ (5)
where $\mathcal{A}_{\text{text}}$ is the emission of free-form natural language tokens via autoregressive decoding; $\mathcal{A}_{\text{action}}$ is a set of structured, non-linguistic actions delimited in the output stream by special tokens <action_start> and <action_end>. These actions may invoke external tools (e.g., call("search", "Einstein")) or interact with an environment (e.g., move("north")), depending on task requirements.
The two subspaces differ in both semantics and functional role: $\mathcal{A}_{\text{text}}$ produces communicative content intended for human or machine interpretation without directly altering the external state, whereas $\mathcal{A}_{\text{action}}$ issues executable commands that either (i) acquire new information through tool calls, or (ii) modify the state of a physical or simulated environment. This separation allows a unified policy to govern both natural language generation and operational decision-making within a single reinforcement learning framework.

### v4

@@ Abstract
The emergence of agentic reinforcement learning (Agentic RL) marks a paradigm shift from conventional reinforcement learning applied to large language models (LLM RL), reframing LLMs from passive sequence generators into autonomous, decision-making agents embedded in complex, dynamic worlds. This survey formalizes this conceptual shift by contrasting the degenerate single-step Markov Decision Processes (MDPs) of LLM RL with the temporally extended Partially Observable Markov Decision Processes (POMDPs) that define Agentic RL. Building on this foundation, we propose a comprehensive twofold taxonomy: one organized around core agentic capabilities, including planning, tool use, memory, reasoning, self-improvement, and perception, and the other around their applications across diverse task domains. Central to our thesis is that reinforcement learning serves as the critical mechanism for transforming these capabilities from static, heuristic modules into adaptive, robust agentic behavior. To support and accelerate future research, we consolidate the landscape of open-source environments, benchmarks, and frameworks into a practical compendium. By synthesizing over five hundred recent works, this survey charts the contours of this rapidly evolving field and highlights the opportunities and challenges that will shape the development of scalable, general-purpose AI agents.
@@ 1 Introduction
The rapid convergence of large language models (LLMs) and reinforcement learning (RL) has precipitated a fundamental transformation in how language models are conceived, trained, and deployed. Early LLM RL paradigms largely treated these models as static conditional generators, optimized to produce single-turn outputs aligned with human preferences or benchmark scores. While successful for alignment and instruction-following, such approaches overlook the broader spectrum of sequential decision-making that underpins realistic, interactive settings. These limitations have prompted a shift in perspective: rather than viewing LLMs as passive text emitters, recent developments increasingly frame them as Agents, i.e., autonomous decision-makers capable of perceiving, reasoning, planning, invoking tools, maintaining memory, and adapting strategies over extended horizons in partially observable, dynamic environments. We define this emerging paradigm as Agentic Reinforcement Learning (Agentic RL). To more clearly delineate the distinction between the concept of Agentic RL studied in this work and conventional RL approaches, we provide the following definition:
@@
In Section 2, we present a more formal, symbolically grounded distinction between Agentic RL and conventional RL. Prior research relevant to Agentic RL can be broadly grouped into two complementary threads: Synergy between RL and LLMs and LLM Agents, detailed as follows:
Synergy between RL and LLMs
The second line of research investigates how reinforcement learning algorithms are applied to improve or align LLMs. A primary branch, RL for training LLMs, leverages on-policy (e.g., proximal policy optimization (PPO) (Schulman et al., 2017) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024b)) and off-policy (e.g., actor–critic, Q-learning (Mnih et al., 2013)) methods to enhance capabilities such as instruction-following, ethical alignment, and code generation (Srivastava & Aggarwal, 2025; Wang et al., 2025m; 2024c). A complementary direction, LLMs for RL, examines the deployment of LLMs as planners, reward designers, goal generators, or information processors to improve sample efficiency, generalization, and multi-task planning in control environments, with systematic taxonomies provided by (Cao et al., 2025c). RL has also been integrated throughout the LLM lifecycle: from data generation (Guo et al., 2025b; Wan et al., 2025a) and pretraining (Dong et al., 2025a) to post-training and inference (Chow et al., 2025), as surveyed by (Guo & Wang, 2025). The most prominent branch here is post-training alignment, notably Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017), along with extensions such as Reinforcement Learning from AI Feedback (RLAIF) (Bai et al., 2022) and Direct Preference Optimization (DPO) (Rafailov et al., 2023; Wang et al., 2024j; Xiao et al., 2024; Liu et al., 2025k; Srivastava & Aggarwal, 2025)
@@
LLM-based agents represent an emerging paradigm in which LLMs act as autonomous or semi-autonomous decision-making entities (Wang et al., 2025d; Li et al., 2025r), capable of reasoning, planning, and executing actions in pursuit of complex goals. Recent surveys have sought to map this landscape from complementary perspectives. Luo et al. (2025a) propose a methodology-centered taxonomy that connects architectural foundations, collaboration mechanisms, and evolutionary pathways, while Plaat et al. (2025) emphasizes the core capabilities of reasoning, acting, and interacting as defining features of agentic LLMs. Tool use, encompassing retrieval-augmented generation (RAG) and API utilization, is a central paradigm, extensively discussed in Li (2025) and further conceptualized by Wang et al. (2024k). Planning and reasoning strategies form another pillar, with surveys such as Masterman et al. (2024) and Kumar et al. (2025) highlighting common design patterns like plan-execute-reflect loops, while Tao et al. (2024) extend this to self-evolution, where agents iteratively refine knowledge and strategies without substantial human intervention. Other directions explore collaborative, cross-modal, and embodied settings, from multi-agent systems (Aratchige & Ilmini, 2025) to multimodal integration (Durante et al., 2024), and brain-inspired architectures with memory and perception (Liu et al., 2025a).
@@
The recent surge in research on LLM agents and RL-enhanced LLMs reflects two complementary perspectives: one explores what large language models can do as the core of autonomous agents, while the other focuses on how reinforcement learning can optimize their behavior. However, despite the breadth of existing work, a unified treatment of Agentic RL, which conceptualizes LLMs as policy-optimized agents embedded in sequential decision processes, remains lacking. Current studies often examine isolated capabilities, domains, or custom environments, with inconsistent terminology and evaluation protocols, making systematic comparison and cross-domain generalization difficult. To bridge this gap, we present a coherent synthesis that connects theoretical foundations with algorithmic approaches and practical systems. We formalize Agentic RL through Markov decision processes (MDPs) and partially observable Markov decision processes (POMDPs) abstractions to distinguish it from classical LLM RL paradigms, and introduce a capability-centered taxonomy that includes planning, tool use, memory, reasoning, reflection (self-improvement), and interaction as RL-optimizable components. Furthermore, we consolidate representative tasks, environments, frameworks, and benchmarks that support agentic LLM training and evaluation, and conclude by discussing open challenges and outlining promising future directions for scalable, general-purpose agentic intelligence. Overall, we aim to further clarify the research scope of this survey:
Primary focus: how RL empowers LLM-based agents (or LLMs with agentic characteristics) in dynamic environments Out of scope (though occasionally mentioned): RL for human value alignment (e.g., RL for harmful query refusal); traditional RL algorithms that are not LLM-based (e.g., MARL (Huh & Mohapatra, 2024)); RL for boosting pure LLM performance on static benchmarks.
@@
This survey is organized to progressively build a unified understanding of Agentic RL from conceptual foundations to practical implementations. Section 2 formalizes the paradigm shift to Agentic RL through an MDP/POMDP lens. Section 3 examines Agentic RL from the capability perspective, categorizing key modules such as planning, reasoning, tool use, memory, self-improvement, perception, and others. Section 6 explores applications across domains, including search, GUI navigation, code generation, mathematical reasoning, and multi-agent systems. Section 5 consolidates open-source environments and RL frameworks that underpin experimentation and benchmarking. Section 6 discusses open challenges and future directions towards scalable, adaptive, and reliable agentic intelligence, and Section 7 concludes the survey. The overall structure is also illustrated in Figure 1.
{forest}
@@ 2 Preliminary: From LLM RL to Agentic RL
LLMs are initially pre-trained using behavior cloning, which applies maximum likelihood estimation (MLE) to static datasets such as web-scraped text corpora. Subsequent post-training methods enhance capabilities and align outputs with human preferences—transforming them beyond generic web-data replicators. A common technique is supervised fine-tuning (SFT), where models are refined on human-generated (prompt, response) demonstrations. However, procuring sufficient high-quality SFT data remains challenging (Maosongcao et al., 2025; Szep et al., 2025; Han et al., 2025). Reinforcement fine-tuning (RFT) offers an alternative by optimizing models through reward functions, circumventing dependence on behavioral demonstrations.
In early RFT research, the core objective is to optimize LLMs through human feedback (Christiano et al., 2017; Ouyang et al., 2022) or data preferences (Rafailov et al., 2023), aligning them with human preferences (RLHF) or directly with data preferences (as in DPO). [^1] Although DPO is another form of optimization objective in RLHF, its complexity is optimized from the perspective of the training process, so it is necessary to distinguish between pure RLHF and DPO. This preference-based RFT (PBRFT) primarily involves learning reward model optimization for LLMs on a fixed preference dataset, or directly implementing it using data preferences. With the release of LLMs such as OpenAI o1 (OpenAI et al., 2024) and DeepSeek-R1 (DeepSeek-AI et al., 2025) that possess reasoning capabilities, their improved performance and cross-domain generalization have garnered widespread attention. With the release of models like OpenAI o3 (OpenAI Team, 2025), which possess both self-evolving reasoning capabilities and support for tool use, researchers are beginning to contemplate how to deeply integrate LLMs with downstream tasks through reinforcement learning methods. Subsequently, researchers have shifted their focus from PBRFT, aimed at optimizing fixed preference datasets, to agentic reinforcement learning tailored for specific tasks and dynamic environments.
In this section, we provide a formalization of the paradigm shift from PBRFT to the emerging framework of agentic reinforcement learning (Agentic RL). While both approaches leverage RL techniques to improve LLMs’ performance, they fundamentally differ in their underlying assumptions, task structure, and decision-making granularity. Figure 2 illustrates the paradigm shift from LLM RL to Agentic RL.
Figure 2: Paradigm shift from LLM RL to Agentic RL. We draw inspiration from (Kumar et al., 2025). The fan-shaped design reflects the outward growth of the RL formulation—from traditional RL (inner), to LLM RL, to full Agentic RL (outer). Color-coded regions represent: red = features specific to LLM RL; teal = features required for Agentic RL; purple = existing Agentic RL implementations. Arrows point outward to indicate increasing interaction breadth (tool use, web browsing, dynamic environments) as one moves toward more agentic settings.
@@ 2.1 Markov Decision Processes
The Markov decision process (MDP) for the RL fine-tuning process can be formalized as a seven-element tuple $\langle\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{P},\mathcal{R},T,\gamma\rangle$, where $\mathcal{S}$ represents the state space and $\mathcal{O}$ is the observation space of the agent. $\mathcal{A}$ denotes the action space. $\mathcal{R}$ is defined as the reward function, $\mathcal{P}$ encapsulates the state transition probabilities, $T$ signifies the task horizon, and $\gamma$ is the discount factor. By casting both preference-based RFT and Agentic RL as MDPs or POMDPs, we clarify the theoretical implications of treating LLMs either as static sequence generators or as interactive, decision-capable agents embedded within dynamic environments.
@@
$\langle\mathcal{S}_{\text{trad}},\mathcal{A}_{\text{trad}},\mathcal{P}_{\text{trad}},\mathcal{R}_{\text{trad}},T=1,\gamma=1\rangle.$ (1)
@@
$\langle\mathcal{S}_{\text{agent}},\mathcal{A}_{\text{agent}},\mathcal{P}_{\text{agent}},\mathcal{R}_{\text{agent}},\gamma,\mathcal{O}\rangle.$ (2)
@@
$\mathcal{A}$: Action space Pure text sequences. $\mathcal{A}_{\text{text}}\cup\mathcal{A}_{\text{action}}$.
$\mathcal{P}$: Transition Deterministic transition to the terminal state. Dynamic transition function $P(s_{t+1}\mid s_{t},a_{t})$.
$\mathcal{R}$: Reward Single scalar $r(a)$. Step‑wise $R(s_{t},a_{t})$; combines sparse task and dense sub‑rewards.
@@ 2.2 Environment State
The LLM agent acts over multiple time-steps in a POMDP. Let $s_{t}\in\mathcal{S}_{\text{agent}}$ denote the full world state and the LLM agent gets observation $O_{t}$ based on the current state $o_{t}=\mathcal{O}(s_{t})$. The LLM agent chooses an action $a_{t}$ based on the current observation $o_{t}$, and the state evolves over time:
$s_{t+1}\sim P(s_{t+1}\mid s_{t},a_{t}).$ (4)
@@ 2.3 Action Space
In the Agentic RL setting, the LLM’s action space comprises two distinct subspaces:
$\mathcal{A}_{\text{agent}}=\mathcal{A}_{\text{text}}\cup\mathcal{A}_{\text{action}}.$ (5)
Here, $\mathcal{A}_{\text{text}}$ denotes the space of free-form natural language tokens emitted via autoregressive decoding, while $\mathcal{A}_{\text{action}}$ denotes the space of abstract, non-linguistic actions, which is usually delimited in the output stream by special tokens <action_start> and <action_end>. These actions may invoke external tools (e.g., call("search", "Einstein")) or interact with an environment (e.g., move("north")), depending on task requirements.
Notably, $\mathcal{A}_{\text{action}}$ is recursively constructed, such that an element $a\in\mathcal{A}_{\text{action}}$ may itself represent a sequence $(a_{1},\dots,a_{k})$ of primitive actions, thus unifying primitive and composite actions within the same space.
Formally, the two subspaces differ in semantics and functional role: $\mathcal{A}_{\text{text}}$ defines the space of outputs intended for human or machine interpretation without directly altering the external state, whereas $\mathcal{A}_{\text{action}}$ defines the space of environment-interactive behaviors that either (i) acquire new information through tool invocations, or (ii) modify the state of a physical or simulated environment. This distinction enables a unified policy jointly to model language generation and environment interaction within the same RL formulation.